The two ME alternatives shown are under consideration for facility improvements in a company in Abu Dhabi. Determine which one should be selected based on a B/C analysis. Assume an interest rate of 10% per year and a 5-year study period.

                                                                                                           Alternative X Alternative Y
First costs, AED 40,000 90,000
Annual M&O costs, AED per year 50,000 20,000
Benefits, AED per year 120,000 150,000
Disbenefits, AED per year 30,000 10,000



Match the closest correct answers for the below questions:

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

What is the total annual cost of Alternative X?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

What is the total annual cost of Alternative Y?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

What is the B/C ratio of Alternative X?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

What is the B/ C ratio of Alternative Y?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

What is the incremental B/C (ΔB/C) ratio between the two alternatives, X and Y?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.       K.   

Which alternative, if any, should be chosen?

A.

[AED 43,742]

B.

[Alternative Y]

C.

[DO NOTHING (DN)]

D.

[-0.59]

E.

[1.49]

F.

[3.2]

G.

[AED 60,552]

H.

[AED 23,742]

I.

[AED 10,552]

J.

[-2.97]

K.

[Alternative X]
Answer #1

Answer a)

Initial Cost of Alternate X AED 40,000

Annual cost is given by, PV= P * (1- (1+r)^-n) / r , P= to be calculated, r = 10%, n= 5 years, PV = AED 40000

Hence 40000 = P * (1-(1+10%)^-5 /10%

40000 = P * (1-(1.1)^-5 /0.1

40000 = P * (1-0.6209) /0.1

40000 = P * 0.3791/0.1

P= 40000 *0.1 / 0.3791

P =AED 10551.31

Hence annual cost is AED 10551.31, in addition AED 50000 is spent on annual O&M. Hence total annual cost is 10551.31+50000 = AED 60551.31

Hence option G is correct

Answer b

Initial Cost of Alternate Y AED 90,000

Annual cost is given by, PV= P * (1- (1+r)^-n) / r , P= to be calculated, r = 10%, n= 5 years, PV = AED 90000

Hence 000000 = P * (1-(1+10%)^-5 /10%

90000 = P * (1-(1.1)^-5 /0.1

90000 = P * (1-0.6209) /0.1

90000 = P * 0.3791/0.1

P= 90000 *0.1 / 0.3791

P =AED 23740.44

Hence annual cost is AED 123740.44, in addition AED 20000 is spent on annual O&M. Hence total annual cost is 23740.44+20000 = AED 43740.44

Hence option A is correct

Answer c)

Total Benefit from Alternate X = AED 120,000

Net Benefit = Benefit - Disbenefit = 120000-30000 = AED 90000

Total Cost of Alternate X = AED 60551.31

Benefit/ Cost Ratio = 90000/60551.31 = 1.486

Hence option E is correct

Answer d)

Total Benefit from Alternate Y = AED 150,000

Net Benefit = Benefit - Disbenefit = 150000-10000 = AED 140000

Total Cost of Alternate Y = AED 43740.44

Benefit/ Cost Ratio = 140000/43740.44 = 3.2

Hence option F is correct

Answer e)

Incremental Benefit from X to Y = Benefit Y - Benefit X = 140000 - 90000 = 50000

Incremental Cost from X to Y = Cost Y - Cost X = 43740.44 - 60551.31 = -16810.87

Incremental Benefit/ Cost = 50000/ (-16810.87) = -2.97

Hence option J is correct

Answer f)

Since Alternate Y has higher Benefit Cost ratio Alternate Y should be chosen

Hence option B is correct